# Assortment Optimization for Patient-Provider Matching

### v1

@@ Abstract
Rising provider turnover forces healthcare administrators to frequently rematch patients to available providers, which can be cumbersome and labor-intensive. To reduce the burden of rematching, we study algorithms for matching patients and providers through assortment optimization. We develop a patient-provider matching model in which we simultaneously offer each patient a menu of providers, and patients subsequently respond and select providers. By offering assortments upfront, administrators can balance logistical ease and patient autonomy. We study policies for assortment optimization and characterize their performance under different problem settings. We demonstrate that the selection of assortment policy is highly dependent on problem specifics and, in particular, on a patient’s willingness to match and the ratio between patients and providers . On real-world data, we show that our best policy can improve match quality by 13% over a greedy solution by tailoring assortment sizes based on patient characteristics. We conclude with recommendations for running a real-world patient-provider matching system inspired by our results.
@@ 1 Introduction
Primary care providers are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022). Patients rely on primary care providers for routine checkups and referrals to specialists. Moreover, care continuity can instill trust and improve medication uptake rates and patient health (Wu et al., 2022).
@@
Unfortunately, high provider turnover rates frequently lead to patients without an assigned provider (Reddy et al., 2015). Provider turnover disrupts patient care and leads to worse care (Reddy et al., 2015). In principle, healthcare administrators reassign unmatched patients to other providers; however, in practice, the process takes months due to provider scarcity and the logistical burden of rematching and coordinating patient matches (Hedden et al., 2021). While many patients find their new provider quickly, others have to wait years to find a new provider due to large numbers of patients, high turnover rates, and provider scarcity (Hedden et al., 2021; Shanafelt et al., 2012).
@@
Algorithms that automatically match patients and providers can reduce logistical hassle but require balancing patient autonomy and system-wide utility. For example, while automatically assigning each patient to a provider would decrease wait times, it also reduces patient autonomy because patients cannot select their provider (Entwistle et al., 2010; Gaynor et al., 2016). In contrast, offering patients full autonomy could lead to suboptimal matches; for example, patients who rematch quickly might prevent better matches for late-responding patients. In addition, granting patients full independence can be overwhelming and delay decisions (Bate and Robert, 2005). We propose to match patients and providers through an assortment framework (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013). Assortments are menus of options given to customers who select an option from such a menu (Bertsimas and Mišic, 2015). Assortments are used in e-commerce settings where platforms give customers a suggested list of products and have them select from this list (Tian et al., 2024). In an assortment setting, each patient receives a menu of providers and selects a provider from such a menu. These curated offer sets allow administrators to balance patient autonomy and system-wide match utility.
We devise policies for offering assortments in a patient-provider matching scenario. To do so, we first develop a model of patient-provider matching then construct a set of policies for offering assortments. We characterize how policy performance depends on problem characteristics such as patient-to-provider ratio, and we empirically detail the regimes under which each of our policies performs best. We then evaluate our policies in a real-world scenario constructed on top of a Medicare dataset to assess the performance of our policies in a real-world example.  [^1] We include all code at https://github.com/naveenr414/patient-provider
Our key question is the following: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and quality?
@@ 1.1 Contributions
We develop a model of patient-provider matching using assortment offerings and analyze its impact on match rate and match quality. We characterize well-performing policies when varying properties such as the patient/provider ratio. We then empirically analyze these policies in a synthetic and real-world dataset built using a Medicare dataset. We use the insights from these studies to propose recommendations for designing patient-provider matching systems in practice.
Developing a Model of Patient-Provider Matching
We develop a model of patient-provider matching where administrators offer assortments upfront, then patients sequentially respond in a random order and select providers one by one. We do so because it captures two critical elements: 1) patients have autonomy in selecting providers and 2) administrators have some control over the matching process by varying assortments offered to patients. We model patient decisions through a choice model that captures how the set of providers offered impacts the final provider chosen. The selection of choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model. Finally, we capture patient heterogeneity through a match quality matrix, which captures differences in match quality between pairs of patients and providers. Match quality can encompass geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988). Patients tend to match with higher quality providers, though the exact match probabilities are dictated by their choice model.
Constructing Assortment Policies
We characterize the optimal policy in the single-provider scenario and use these insights to construct a set of policies for assortment optimization. We first construct a baseline policy, greedy, which offers every provider to every patient. We demonstrate that the greedy policy can do arbitrarily poorly, so we build upon this by developing three new policies that adaptively vary assortment sizes: pairwise, group-based, and gradient descent. The pairwise policy offers each patient at most one provider, and we demonstrate an instance-dependent approximation guarantee for this policy. We then develop the group-based policy by augmenting the pairwise policy so we offer assortments to groups of patients. We demonstrate that such group-based maintains the match rate of the pairwise policy while potentially improving the match quality. For the gradient descent policy, we construct an objective function that provably lower bounds the match quality, and we use gradient descent to optimize this quantity.
Characterizing Policy Performance
We analyze the performance of our policies in a synthetic setting and empirically characterize the conditions under which each policy performs best. We demonstrate that the best-performing policy depends on two factors: the match probability and the ratio of patients to providers. When patients outnumber providers, as is common in many real-world health systems, the gradient descent policy performs best because the objective function is closer to the lower bound. When patient and provider counts are balanced, we show that the greedy policy achieves a high match quality when patients have low match probabilities, while the group-based policy achieves a high match quality when patients have high match probabilities.
We evaluate policies in a real-world scenario constructed via Medicare data (Medicare & Medicaid, 2025) and we show here that the gradient descent policy performs best. We extend this scenario to real-world considerations such as fairness and cognitive overload.
Contribution Summary
@@
Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) develop a model of patient-provider matching under assortments, ii) characterize the performance of assortment policies, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) make recommendations for the implementation of a real-world patient-provider matching system based on the results of our theoretical and empirical analysis.
@@
Figure 1: A) In this work, we match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers. C) Patients then respond sequentially and either select a provider from their assortments or abstain. D) This results in matches between patients and providers.
@@ 1.2 Related Work
Matching Patients and Providers Care continuity is critical for patient health, as it allows for better patient communication, lowers operating costs, and improves provider teamwork (Plomondon et al., 2007). Despite this, provider turnover rates are high, with the average healthcare system reporting a 7% turnover rate year-to-year (Plomondon et al., 2007). Moreover, such issues are exacerbated recently due to high rates of provider burnout (Shanafelt et al., 2012) and the Covid-19 pandemic (Shanafelt et al., 2022). Once a provider drops out, it is difficult for patients to find a new one; while 54% of patients find their new provider within 12 months, 6% fail to find a new provider even after 36 months (Hedden et al., 2021). High levels of provider turnover can worsen primary care because patients lack access to routine checkups or reminders to take medication (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019).
While provider turnover impacts patient satisfaction, quantifying “match quality” is hard. Patients care mainly about geographic proximity (Dunlea and Lenert, 2015), with concordance along race (Greenwood et al., 2020), gender (Greenwood et al., 2018), and language (Manson, 1988) playing a secondary role. Communication style alignment can improve match quality (Street et al., 2008). These factors build trust and lead patients to more truthfully report symptoms (Persky et al., 2013).
Algorithms for Patient-Provider Matching Existing work in patient-provider matching uses either a one-shot matching framework or a genetic programming-based approach. Within the one-shot framework, linear programs can manage provider panel sizes (Harrington et al., 2021), the deferred acceptance framework can perform two-stage matching (Chen et al., 2020), and scheduling algorithms are used to model optimal rates of on-demand scheduling (Qu et al., 2007).
@@
A second approach to patient-provider matching builds matches through genetic programming. Genetic programming algorithms learn matches over time through a fitness function. Within this framework, one line of work uses genetic programming to find fair matches between patients and providers (Chen et al., 2020), while another approach uses genetic programming to balance workloads between providers (Zhu et al., 2023). We build on these lines of work by incorporating patient autonomy into the process, which motivates the need for an assortment-based approach.
@@
Assortment Optimization Our work analyzes patient-provider matching through assortment optimization, where retailers offer a set of menus to customers. In assortment optimization, retailers construct assortments and customers make decisions based on offered options (Bertsimas and Mišic, 2015). Retailers first construct models of customer decisions, such as the multinomial logit model (Davis et al., 2013) and nested logit models (Alfandari et al., 2021), both of which use logit probabilities to model customer decisions. Assortments are employed in domains including school choice (Shi, 2016), dating markets (Rios and Torrico, 2023), and e-commerce (Tian et al., 2024).
Assortment optimization algorithms come in offline and online variants. In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015). In online assortment optimization, customers arrive sequentially, and options are offered for each (Aouad and Saban, 2023). Our setting resides between offline and online, as we offer assortments offline, but patients respond online.
Matching under Random Ordering Patient-provider matching is challenging from an optimization perspective because patients make decisions in a random order. Randomized response orders are found in variants of online selection problems such as secretary problems (Freeman, 1983) and online learning (Orabona, 2019). One key difference is the lack of flexibility in our scenario; changing assortments online for the patient-provider matching setting would involve considerable logistical hassle, so we offer assortments in a one-shot manner.
@@ 2.1 Formal Model of Patient-Provider Matching
We focus on matching $N$ patients to $M$ providers. We present each patient $i$ with an assortment $\mathbf{X}_{i}\in\{0,1\}^{M}$ upfront. Here, $X_{i,j}$ denotes whether patient $i$ has provider $j$ in their assortment. Patients then select providers from their assortment sequentially by responding in an order $\sigma$, and can abstain from provider selection. Here, $\sigma_{1}\in[N]$ represents the first patient in the order. We assume patient response times are i.i.d. and exponential; prior work demonstrates this can model real-world wait times (Newell, 2013). Under this assumption, $\sigma$ is a random permutation of $[N]$.
Each patient selects a provider from their assortment according to a choice model $f_{i}:\{0,1\}^{M}\rightarrow\{0,1\}^{M}$. The choice model takes as input a 0-1 vector $\mathbf{z}$ and outputs a random 0-1 vector representing the provider selected by patient $i$. Formally, let $\mathbf{y}^{(t)}$ represent the set of unselected providers at time $t$. Then, patient $\sigma_{t}$ selects providers according to $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})$, where $\odot$ is the element-wise product ($\mathbf{a}\odot\mathbf{b}=\mathbf{c}\rightarrow a_{k}b_{k}=c_{k}\forall k$). $\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}$ represents the set of providers that are on patient $\sigma_{t}$’s assortment and unselected by previous patients. If $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})=\mathbf{0}$, then no provider is selected, while if $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})_{j}=1$, then provider $j$ is selected. We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$
Patients are more likely to select providers with a higher match quality, $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ when selecting provider $j$. Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs. We assume administrators know $\theta_{i,j}$. Prior work demonstrates providers are able to successfully identify new high-match quality providers for patients (Lohr et al., 2013), which implies that we can predict match quality from patient data.
We present examples of choice models below:
@@
    Uniform - Patients select their most preferred provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
@@
    Threshold - Patients follow the uniform choice model but only select providers with match quality above some threshold $\alpha$. If match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance. Formally:
@@
    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option $\gamma$ which represents selecting no provider. Under the MNL choice model (Train, 2009), this is:
@@
We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$, given the match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
@@
    Match Quality measures $\theta$ across selected patient-provider pairs. Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (4)
@@
    Match Rate is the proportion of patients who are matched. Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\rVert_{1}]$ (5)
Example 1.
We quantify the example in Figure 1. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be

### v2

@@ Abstract
Rising provider turnover results in frequently needing to rematch patients with available providers. However, the rematching process is cumbersome for both patients and health systems, resulting in labor-intensive and ad hoc reassignments. We propose a novel patient-provider matching approach to address this issue by offering patients limited provider menus. The goal is to maximize match quality across the system while preserving patient choice. We frame this as a novel variant of assortment optimization, where patient-specific provider menus are offered upfront, and patients respond in a random sequence to make their selections. This hybrid offline-online setting is understudied in previous literature and captures system dynamics across various domains. We first demonstrate that a greedy baseline policy–which offers all providers to all patients–can maximize the match rate but lead to low-quality matches. Based on this, we construct a set of policies and demonstrate that the best policy depends on problem specifics, such as a patient’s willingness to match and the ratio of patients to providers. On real-world data, our proposed policy improves average match quality by 13% over a greedy solution by tailoring assortments based on patient characteristics. Our analysis reveals a tradeoff between menu size and system-wide match quality, highlighting the value of balancing patient choice with centralized planning.
@@ 1 Introduction
Primary care providers (PCPs) are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022). Patients rely on PCPs for routine checkups and referrals to specialists. PCPs serve as a regular touchpoint for patients and provide care continuity, which has been shown to instill trust and improve medication uptake rates and patient health (Wu et al., 2022). Furthermore, high-quality patient-provider matches compound the benefit of primary care, where the notion of a high-quality match is heterogeneous across patients and can include factors such as geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988). As one such example, Alsan et al. (2019) found that access to race-concordant doctors could reduce cardiovascular disease by 19% (Alsan et al., 2019).
@@
Unfortunately, high provider turnover rates frequently lead patients to lose their PCP, which disrupts patient care and can lead to worse health outcomes (Reddy et al., 2015). In principle, healthcare administrators reassign unmatched patients to other providers. However, in practice, the process takes months due to provider scarcity and the administrative burden of patient coordination (Hedden et al., 2021).
@@
As an alternative to manual reassignments, algorithmically matching patients and providers can reduce logistical hassle. However, this approach requires carefully balancing patient autonomy and system-wide utility. While automatically assigning each patient to a provider would decrease wait times, it also eliminates patient choice (Entwistle et al., 2010; Gaynor et al., 2016). In contrast, offering patients full autonomy to choose between all available providers could lead to suboptimal matches when considering system-wide match quality; for example, patients who choose quickly might prevent better matches for late-responding patients. In addition, granting patients full independence can be overwhelming and lead to decision delay (Bate and Robert, 2005).
In this work, we formulate the patient-provider matching problem as a variant of assortment optimization (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013); we illustrate this problem in Figure 1. Assortments are menus of options given to customers, who then make selections from this menu (Bertsimas and Mišic, 2015). This framework is commonly seen on e-commerce platforms (Tian et al., 2024). By offering limited provider menus, patients have the autonomy to select their preferred provider, while healthcare administrators can curate these assortments to maximize system-wide metrics.
Our work studies a unique variant of assortment optimization with offline assortment and online response: assortments are created and offered simultaneously, and patients then respond sequentially in a random order to make provider selections. This variant is especially interesting because it captures the logistical complexities of the patient-provider matching problem. Dynamically varying assortments online are logistically taxing for healthcare administrators, yet it is difficult to control the patient response order, necessitating a combination of offline and online modeling. Moreover, this setting generalizes beyond patient-provider matching; similar situations can be found in other two-sided matching markets such as food delivery, where companies may propose a set of order options (menus) to drivers who respond in an online fashion, and where there is a tradeoff between driver autonomy and system utility (i.e., cost).
Our contributions answer this key question: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and match quality? Our results show that limiting assortment sizes and offering carefully tailored menus enables a tradeoff between match rate and match quality, underscoring the balance between patient autonomy and centralized planning.
@@ 1.1 Contributions
Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) model patient-provider matching using a hybrid offline-online variant of assortment optimization, ii) characterize the performance of assortment policies theoretically and empirically, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) demonstrate that carefully tailoring assortment menus allows us to tradeoff between match rate and match quality.
Patient-Provider Matching as Assortment Optimization
We develop a model of patient-provider matching where administrators offer assortments upfront, and then patients respond sequentially in a random order. We do so to capture two critical elements: 1) patients have autonomy in selecting providers, and 2) the system has some control over the matching process in selecting assortments for each patient. Additionally, such a model can be broadly applicable to two-sided marketplaces where one side (users) selects online from an assortment, yet dynamic variation of the assortment is infeasible. We model patient decisions through a choice model that captures how the set of offered providers impacts the resultant matches. The selection of the choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model. Finally, we capture patient heterogeneity through a match quality matrix, which captures differences in match quality between pairs of patients and providers.
Characterizing Assortment Policies
We study various policies for assortment optimization and characterize their performance theoretically and empirically. We quantify policy performance through two metrics: match rate and match quality. We first propose a baseline greedy policy that offers all providers to all patients, then demonstrate that this policy can perform arbitrarily poorly for match quality. We then build upon this by developing new policies that adaptively vary assortment sizes. We demonstrate that tailored assortment offerings can improve match quality compared to baseline greedy solutions. We empirically characterize the best policy as a function of model parameters such as the ratio of patients to providers and the underlying model assumptions.  [^1] We include all code at https://github.com/naveenr414/patient-provider/
Real-World Matching System Insights
We apply our policies in a semi-synthetic simulation based on a real-world healthcare system in Connecticut. We demonstrate that the best-performing policy here outperforms the greedy baseline by 13% for match quality. We extend this to consider metrics beyond match quality and match rate, such as fairness and regret. These findings yield actionable recommendations and highlight the tensions between various metrics when designing real-world patient-provider matching systems. Our key takeaway is that the best set of assortments offered depends on problem specifics, including the ratio of patients to providers and the willingness of patients to match. As patients are more willing to match, assortments should be smaller and more tailored to maximize match quality.
@@
Figure 1: Overview of patient-provider matching model. A) We match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers. C) Patients then respond sequentially in a random order and either select a provider from their assortments or abstain. D) This results in matches between patients and providers.
@@ 1.2 Related Work
Matching Patients and Providers Care continuity is critical for patient health, as it allows for better patient communication, lowers operating costs, and improves provider teamwork (Plomondon et al., 2007). Despite this, provider turnover rates are high, with the average healthcare system reporting a 7% turnover rate year-to-year (Plomondon et al., 2007). Moreover, such issues are exacerbated recently due to high rates of provider burnout (Shanafelt et al., 2012) and the Covid-19 pandemic (Shanafelt et al., 2022). When a provider leaves, it is difficult for patients to find a new one; only 54% of patients find their new provider within one year, and 6% fail to find a new provider even after three years (Hedden et al., 2021). High levels of provider turnover can worsen primary care because patients lack access to routine checkups which can reduce accountability, leading to consequences such as lower medication adherence (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019). Provider turnover itself impacts patient satisfaction and health outcomes, but the quality of the new match is also consequential in both dimensions. Nevertheless, it is difficult to quantify what constitutes a “high-quality” match. Patients care mainly about geographic proximity (Dunlea and Lenert, 2015), with concordance along race (Greenwood et al., 2020), gender (Greenwood et al., 2018), and language (Manson, 1988) playing a secondary role. Communication style alignment can improve match quality (Street et al., 2008). These factors build trust which translates to better care, such as leading patients to more truthfully report symptoms (Persky et al., 2013).
@@
Algorithms for Patient-Provider Matching Existing work in patient-provider matching typically adopts either a one-shot matching framework or a genetic programming-based approach. Within the one-shot framework, linear programs are used to manage provider panel sizes (Harrington et al., 2021); deferred acceptance frameworks enable two-stage matching (Chen et al., 2020); and scheduling algorithms model optimal rates of on-demand scheduling (Qu et al., 2007). An alternative approach to patient-provider matching uses genetic programming, which learns matches over time through a fitness function. Within this framework, one line of work uses genetic programming to find fair matches between patients and providers (Chen et al., 2020), while another approach uses genetic programming to balance workloads between providers (Zhu et al., 2023). In contrast to these approaches, we incorporate patient autonomy into the process, which motivates our assortment-based approach.
@@
Assortment Optimization Our work analyzes patient-provider matching through assortment optimization. In this setting, retailers construct assortments for customers, and customers make decisions based on offered options (Bertsimas and Mišic, 2015). Retailers use models to capture customer choice behavior, such as the multinomial logit (Davis et al., 2013) and nested logit (Alfandari et al., 2021), and generate product menus accordingly. Assortments are ubiquitous in e-commerce settings (Tian et al., 2024) but also employed in other domains such as school choice (Shi, 2016) and dating markets (Rios and Torrico, 2023). Both offline and online variants have been studied. In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015), while in online assortment optimization, customers arrive sequentially, and options are dynamically offered for each (Aouad and Saban, 2023). Other variants of assortment optimization include settings where system parameters are unknown (Rusmevichientong and Topaloglu, 2012), are dynamically learned (Caro et al., 2014), or have constraints (Feldman and Topaloglu, 2018). Our setting resides between offline and online, as we offer assortments offline, but patients respond online.
@@ 2.1 Formal Model of Patient-Provider Matching
We introduce the problem of matching $N$ patients to $M$ providers. We present each patient $i$ with an assortment $\mathbf{X}_{i}\in\{0,1\}^{M}$ upfront. Here, $X_{i,j}$ denotes whether patient $i$ has provider $j$ in their assortment. Patients then select providers from their assortment sequentially by responding in an order $\sigma$, and can abstain from provider selection. Here, $\sigma_{1}\in[N]$ represents the first patient in the order. We assume patient response times are i.i.d. so $\sigma$ is a uniformly random permutation of $[N]$; prior work demonstrates this can model real-world wait times (Newell, 2013) and we relax this assumption in Section 4.3.
Patients select providers based on match quality $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ selecting provider $j$. Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs. We assume administrators know $\theta_{i,j}$ because $\theta_{i,j}$ can be estimated through patient surveys; we further discuss this assumption in Section 2.2.
Each patient selects a provider from their assortment according to a choice model $f_{i}:\{0,1\}^{M}\rightarrow\{0,1\}^{M}$. The choice model takes as input a 0-1 vector $\mathbf{z}$ denoting available and offered providers, and it outputs a random 0-1 vector representing the provider selected by patient $i$. Formally, let $\mathbf{y}^{(t)}$ represent the set of unselected providers at time $t$. Then, patient $\sigma_{t}$ selects providers according to $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})$, where $\odot$ is the element-wise product ($\mathbf{a}\odot\mathbf{b}=\mathbf{c}\rightarrow a_{k}b_{k}=c_{k}\forall k$). $\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}$ represents the set of providers that are on patient $\sigma_{t}$’s assortment and unselected by previous patients. If $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})=\mathbf{0}$, then no provider is selected, while if $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})_{j}=1$, then provider $j$ is selected. We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$. We assume that each provider can match with at most one patient as well. We assume this to simplify the presentation of our paper, though we relax this assumption in Section 4.3 and show in Appendix C that this is empirically equivalent to modifying the total number of providers. We present examples of choice models below:
@@
    Uniform - Patients select their most preferred available provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
@@
    Threshold - Patients follow the uniform choice model but only select providers with match quality above some threshold $\alpha$. For example, if match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance. Formally:
@@
    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option valued at $\gamma$ which represents selecting no provider. Under the MNL choice model (Train, 2009), $f$ is:
@@
We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$ from a match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
@@
    Match Rate (MR) is the proportion of patients who are matched. Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
    $\mathrm{MR}(\mathbf{X},\theta,f)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\rVert_{1}]$ (4)
@@
    Match Quality (MQ) measures $\theta$ across selected patient-provider pairs. Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
    $\mathrm{MQ}(\mathbf{X},\theta,f)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (5)
We select these two metrics because they naturally capture the scale and quality of matches, though, in Section 5, we evaluate additional metrics such as fairness and regret.
Example 2.1.
We quantify the example in Figure 1 to demonstrate the metric calculations. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be